OpenStudy (anonymous):

To evaluate the sine, cosine, and tangent of the angle, first find a coterminal angle of the given angle.

OpenStudy (anonymous):

Yes.. this is what you do... but there isn't a problem here.

OpenStudy (anonymous):

i dunno

OpenStudy (anonymous):

what is it that you don't know? :)

OpenStudy (anonymous):

how to find the coterminal angle

OpenStudy (anonymous):

a coterminal angle is an angle that is obtained when you either subtract or add the angle with 360 degrees :)

OpenStudy (anonymous):

example: Let's say we have A = -200 degrees, find the positive and negative coterminal angle In this case, the positive coterminal angle is Coterminal angle = -200 + 360 = 160 degrees Coterminal angle (negative) = -200 -360 = -560 degrees are you understanding it now bit by bit? :)